OpenStudy (anonymous):

(cos Θ − cos Θ)2 + (cos Θ + cos Θ)2

OpenStudy (anonymous):

@mathmale

OpenStudy (mathmale):

JSD: First I'm going to ask you to please type in those exponents properly. The square of x would properly be typed in as x^2. Secondly, please double-check to ensure that you've copied this problem down properly. I ask you to do this because \[(\cos \theta-\cos \theta)=0.\]

OpenStudy (anonymous):

(cos Θ − cos Θ)^2 + (cos Θ + cos Θ)^2

OpenStudy (anonymous):

thats right

OpenStudy (mathmale):

what is the value of cos theta - cos theta?

OpenStudy (mathmale):

sorry I have to excuse myself, I'll be back

OpenStudy (anonymous):

would it be cos^2θ then?

OpenStudy (anonymous):

and alright

OpenStudy (anonymous):

and to answer your question ) correct

OpenStudy (anonymous):

0*

OpenStudy (mathmale):

Still on the phone but will help you as far and as soon as I can. no, cos theta - cos theta would be zero, not (cos theta)^2.

OpenStudy (anonymous):

yah i know i said zero i thought that was the answer

OpenStudy (anonymous):

and nah take your time

OpenStudy (mathmale):

You have two squares. If you've copied the problem down correctly, the first (the left one) is zero. The second is (cos theta + cos theta)^2. Do you agree with that? If so, evaluate it. (Still on phone.)

OpenStudy (mathmale):

Hint: add cos theta to cost theta. Result?

OpenStudy (anonymous):

yah i agree thats whu i thought it was cos^2θ… and it would be cos^2 correct

OpenStudy (anonymous):

?

OpenStudy (anonymous):

oooh wait

OpenStudy (mathmale):

Important that you write this as (2 cos x)^2 or \[(2 \cos \theta )^2~or~2^2\cos ^{2}\theta \]

OpenStudy (anonymous):

alright so same thing

OpenStudy (anonymous):

so when you add ^2 it would be 4cos^2 Θ correct?

OpenStudy (mathmale):

Essentially, yes. You're not "adding ^2"; instead, you are SQUARING 2 cos theta.

OpenStudy (anonymous):

thats what i ment sorry

OpenStudy (mathmale):

I'm picky because this language needs to be precise.

OpenStudy (mathmale):

So, what do you believe is the final answer? Please follow one of the examples of formatting that I've given you.

OpenStudy (anonymous):

I understand and you are right ! thx for the help i understand it a lot better now and thx for walking me through it.

OpenStudy (anonymous):

4cos^2 theta

OpenStudy (mathmale):

i would still prefer that you write that as 4 (cos theta)^2 or \[4 \cos ^{2}theta\] for maximum clarity.

OpenStudy (mathmale):

Any further questions about this problem?